one alloy contains 80% gold (in weight) and another alloy contains 55% gold. how many grams of each alloy should be contained to make a 40kg of an alloy that contains 70% gold?
Answer by mananth(16946)

80% gold (in weight) let this be x kg
another alloy contains 55% gold. this will be 40-x kg
Mix 40kg of an alloy that contains 70% gold?
..
80x+55(40-x)=40*70
80x+2200-55x=2800
25x=2800-2200
25x=600
x=600/25
x= 24 kg ----- 80% gold
balance 16 kgs will be 55% gold alloy
...
m.ananth@hotmail.ca